OpenStudy (anonymous):

Solve for d. -10 = -2 + 1/2 d

OpenStudy (anonymous):

-8.5=d

OpenStudy (anonymous):

WRONG

OpenStudy (anonymous):

O.O

OpenStudy (anonymous):

A. d = -24 B. d = -16 C. d = -6 D. d = -4

OpenStudy (anonymous):

-2+.5= -1.5... O.o is the 1/2 TIMESING the d?

OpenStudy (anonymous):

-8=1/2 d -4=d

OpenStudy (anonymous):

-10 = -2 + 1/2 d

OpenStudy (anonymous):

Wait wait... the 1/2 is being MULTIPLIED by the D?! :O